# TIP-Search: Time-Predictable Inference Scheduling for Market Prediction under Uncertain Load

### v1

@@ I Introduction
LOB Task InputDeadline MonitorModel SelectorModel A(Low Latency)Model B(Balanced)Model C(High Accuracy)Prediction OutputModel Pool
@@ VIII-C Visualization of Scheduling Efficacy
To further elucidate the effectiveness of each strategy, Figure 9 presents a heatmap of task source versus selected model accuracy. It illustrates that TIP-Search adapts to input variability by selecting high-performing models per task, thereby avoiding low-accuracy assignments such as those frequently triggered by the FI-2010 model. In contrast, other strategies exhibit rigid or naive selection behaviors, leading to degraded accuracy on mismatched inputs.
Figure 9: Accuracy heatmap showing model-task pairings across scheduling strategies.

### v2

@@ VIII-C Visualization of Scheduling Efficacy
To provide a comprehensive view of how different strategies assign models to input data sources, Figure 9 presents an accuracy heatmap that combines results from the Fixed, Random, and Source-Only strategies. Each row corresponds to a (strategy, input source) pair, and each column represents a candidate model. The cell values denote the classification accuracy achieved when that model is selected for the given source under the respective strategy.
The heatmap clearly illustrates the limitations of naive or static strategies. For example, the Fixed-FI2010 configuration performs poorly across all models except its source, whereas Random strategies fail to consistently align high-accuracy models with appropriate inputs. Source-Only tends to overfit to its domain-specific models, limiting generalization. These patterns highlight the value of adaptive mechanisms like TIP-Search, which dynamically select models based on real-time performance profiles.
Figure 9: Accuracy heatmap for (strategy, input source) vs. selected model. Values reflect achieved accuracy per configuration.